Could you please help me with this question?: The average (arithmetic mean) of ten positive numbers is 16. The average of the least and the greatest of these numbers is 19. What is the average of the other eight numbers?
Answer by Alan3354(30993)

The average (arithmetic mean) of ten positive numbers is 16. The average of the least and the greatest of these numbers is 19. What is the average of the other eight numbers?
--------------
10*16 = 160 (10 times the avg of 10)
2*19 = 38 (2 times the avg of 2)
----
160 - 38 = 122 (8 times the average of the remaining 8)
122/8 = 15.25
Avg of 8 = 15.25